OpenStudy (anonymous):

Double integral help! I don't think this is solvable because Wolfram Alpha timed out lol. Any help would be appreciated!

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{\sqrt{y}}^{1}\sqrt{x^3 +1}dxdy\]

OpenStudy (anonymous):

have you tried changing your integral? instead of it being dxdy try to change it to dydx. im actually doing this at the moment so ill let you know what i got

OpenStudy (anonymous):

I just tried it and maybe I'm doing it wrong (because we haven't gone over it in class, but it's due) but I got the same problem. I don't know how to do:\[\int\limits_{\sqrt{y}}^{1}\sqrt{x^3+1} dx\]

OpenStudy (anonymous):

is it online or do you have the answer or is it book work?

OpenStudy (anonymous):

try doing it this way \[\int\limits_{0}^{2}\int\limits_{0}^{\sqrt{x}}\sqrt{x^3+1}dydx\]

OpenStudy (anonymous):

I have the answer. It's\[(\frac{ 2 }{ 9 })(2^{3/2}) - (\frac{ 2 }{ 9 })\]

OpenStudy (anonymous):

the answer i got was 3.416009726 but i did it kind of quickly

OpenStudy (anonymous):

Why did you switch the bounds like that? Sorry, I haven't learned this in class yet.

OpenStudy (anonymous):

you can switch the bounds depending on if its dxdy or dydx

OpenStudy (anonymous):

some problems arent possible without switching the bounds

OpenStudy (anonymous):

Well I mean, what did you do to get the bounds as they are in your post?

OpenStudy (anonymous):

its hard to explain over a computer

OpenStudy (anonymous):

okay, thanks for your help! I'll try to google some stuff

OpenStudy (anonymous):

sorry i wasnt much help:/

OpenStudy (anonymous):

No, you were! I didn't know you had to switch the bounds that way. That'll help